Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2025 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 8     Entry time: 31 Jan 2025, 21:24
Author: Chan Kim 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 11.0 keV, up to 352 pulses at 1.1 MHz, 900 uJ.

Optical laser delivery

  • ns illumination and pp lasers

Achievements/Observations

  • Finished with Ice detection program
  • Optimized the beamline alignment for the next experiment
  • Checked the NKB focusing with CRL collimation: The collimated scheme resulted in smaller focus size and significantly higher intensity at focus

Issues

  • PAM mirror mount is not reacheable with the SmarAct motor
ELOG V3.1.4-7c3fd00